I need help solving this one, if anyone could help?
2x-5y=11, 4x+10y=18
: I need help solving this one, if anyone could help?
2x-5y=11, 4x+10y=18

Answer by Alan3354(1931)

I need help solving this one, if anyone could help?
2x-5y=11, 4x+10y=18
---------------
2x- 5y=11 eqn 1
4x+10y=18 eqn 2
4x-10y=22 Multiply eqn 1 by 2
Add the 2 eqns
8x = 40
x = 5
Sub into either eqn
2*5 - 5y=11
10 - 5y = 11
-5y = 1
y = -1/5